OpenStudy (anonymous):

(-3)^4 (-3) * (-3) = 9 * (-3) = 27 * (-3) = 81 Check this one, please (-3)^4 (-3) * (-3) = 9 * (-3) = 27 * (-3) = 81 Check this one, please @Mathematics

OpenStudy (anonymous):

It's 81 (-3) * (-3) = 9 * (-3) = -27 * (-3) = 81

OpenStudy (anonymous):

since the power is even ,the answer must be positive

OpenStudy (anonymous):

Thanks